Dear all,

please find below a short summary of the activities of the 27th related to the AO sys

People at LBTO:G. Brusa

people working in remote:A. Riccardi (VNC+Skype)

March 27th Total # of person: 2

Summary:

The night started reasonably well with ASM set around UT 1:40 until ~0500 UT, after two resets of the ASM followed in less than 15 minutes, they appear to be due to some over-current (TBC). After the second reset the incorrect recovery of the software caused an anomalous condition of operation in which the elevation of the telescope was not correctly reported to the ASM which in turn caused the ASM current thresholds to be incorrectly set and a long list of failed attempts to set the shell. After about two hours of struggling we finally determined the cause of the error and the ASM restarted properly (UT 08:00). One more rest occurred at UT 10:58 and the observations terminated at UT 13:04.

Short log:

UT 01:40 Dome open (including side and rear), telescope horizon pointing, shell at rest position. Wind

4-5 m/s. Cooling flow (9.6 l/min), cooling inlet 0.9C, outlet 2.7C, hub -0.4, CP temp 2.1C, Ref-Body 1.7C.

UT 01:51 Slewing towards object

UT 01:52 setting ASM.

UT 01:54 ASM set at 78 deg elevation, taken the offset for forces and position. First star give a pretty good collimation.

UT 04:28 Seeing still very bad, proceeding with some data acquisition for the pointing model. Tracking new object at 79 deg elevation, wind speed around 2 m/s. Cooling flow (9.3 l/min), cooling inlet -1.7C, outlet 0.6C, hub -4.9, CP temp -1.4C, Ref-Body -2.4C.

UT 05:09 ASM reset itself.

UT 05:12 Recovery using AdSec Control successful. The shell is now in REF state and the ‘classic ear’, i.e. a small are around 7 o’clock and at the outer edge has appeared.

UT 05:18 the ASM was set and is running OK.

UT 05:23? ASM reset itself again.

UT 05:39 Calling Armando in Italy.

(here follows a long list of mirror resets)

UT 08:00 The ASM was finally set! Elevation 65, wind 4 m/s, flo 9.2 l/min, cooling inlet -2.4C, outlet -0.4C, hub -5.0C, CP -1.9C, ref body -2.5C.

UT 09:51 the ASM is working properly. Telescope pointing at 45 deg elevation , with wind speed at 4 m/s. Cooling at flow rate 9.3 l/min, inlet -1.9C, outlet 0.4C, hub -5.7C, CP -1.7C, ref body -2.6 C.

UT 10:58 ASM found REF-ed, started the Recover Fail from AdSec Control.

UT 11:02 ASM correctly re-started and set.

UT 12:55 Several attempts to rest the ASM via the AdOptControl were unsuccessful, reporting NOT PRCESSED, after consultation with Armando used the Rest command on the AdSec Control GUI.

UT 13:04 ASM in rest position with all the GUIs agreeing: AdoptControl saying of AdSec System Process UP and Hardware Status OPERATING, AdSec Control GUI saying AdSec Arbitrator Status: Operating, SHELL RIP, TSS OFF, POWER ON.

Ciao,

Guido

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback